# Modular memristor model with synaptic-like plasticity and volatile memory

### 05 Mar 2026 v1

@@ V Discussion
V.1 Extending the state-of-the-art memristor models
@@
To model synaptic plasticity, we implement a spike-timing-dependent plasticity (STDP) rule inspired by the work of Gerstner et al. [6]. This approach is based on the concept of eligibility traces, a cornerstone of modern theories of learning in the brain. In this framework, plasticity is a two-stage process. First, the near-coincidence of pre- and post-synaptic activity creates a short-lived synaptic "tag" or trace that marks the synapse as "eligible" for modification. Second, this temporary trace is only converted into a lasting weight change if a third, modulatory signal—representing global factors like reward, novelty, or surprise—arrives while the trace is still active [7]. By adopting this mechanism, our model moves beyond simple pairwise STDP to support more powerful and flexible three-factor learning rules. This capability aligns the model with cutting-edge research in reinforcement learning and makes it a compelling component for building truly adaptive neuromorphic systems.
@@
V.2 Physical origin of the 1/t-like memory kernel
As shown in Section IV.2, the volatility module of our model is governed by the hereditary kernel in Eq.˜9 with an exponent $\alpha\approx 0.03$ obtained from the log–log relaxation data in Figure˜2. The corresponding cumulative memory $K(t)=\int_{-\infty}^{t}\ker(\tau)\,d\tau$ decays approximately as $K(t)\sim 1/t$ for large $t$. In practice, the device therefore exhibits effectively $1/t$-like “loss of memory” over time.
@@ V.2 Physical origin of the 1/t-like memory kernel
The viscoelastic analogy used in Section II.1 provides a convenient heuristic explanation of this situation. In linear viscoelasticity, a material with a broad spectrum of relaxation times $\rho(\tau)$ can be represented as a superposition of many Maxwell modes, and a spectrum $\rho(\tau)\propto\tau^{-1-\alpha}$ yields an effective relaxation modulus $G(t)\propto t^{-\alpha}$ [15, 22]. Our choice of $\ker(t)\propto(t+\varepsilon)^{-(\alpha+1)}$ plays the same role: it compactly encodes the presence of many internal relaxation processes without having to resolve them individually at the level of the model.
@@
V.3 Adapting model to polymeric memristor
@@
V.4 Prospective model extensions
@@
VI Conclusions
We have presented a modular and computationally efficient compact model for memristors that exhibits both volatile memory and synaptic-like plasticity. The model’s key innovation lies in the synthesis of several modeling techniques that allow for memristor current-voltage characteristics, synaptic-like plasticity effects and a general description of relaxation dynamics (volatile memristor memory) with saturation. We have discussed how the volatile relaxation dynamics, experimentally observed as a power-law decay, can be described by a viscoelasticity-inspired hereditary integral whose form is physically justified by percolation theory. Furthermore, we have demonstrated that a biologically plausible model of STDP based on eligibility traces can be implemented, endowing the future devices with a potential capacity for advanced learning.
The complete model was shown to be in excellent quantitative agreement with experimental data from polymeric memristors. However, we believe that the modularity and relative simplicity of the proposed model allow its implementation (including validation) in a wide range of memristive-like devices. Therefore, this work provides a unified, mathematically grounded framework for understanding and simulating complex memristive devices, offering a valuable tool for advancing the design of neuromorphic computing systems.
Acknowledgments
@@
D.H. and L.K. conceived and led the project, carried out the theoretical development, implemented the model and simulations, analyzed the data, and wrote the manuscript. S.H.F. and K.K. contributed to the theoretical development and provided critical feedback and revisions to specific sections of the manuscript. A.P., Y.R.P., J.P., and J.V. fabricated the polymeric memristors and performed the experimental measurements. All authors reviewed the manuscript and approved the final version.
@@
Appendix A Numerical values
@@
The first table˜1 lists the kernel exponent $\alpha$ obtained from the analysis presented in Section˜IV.2. The second table˜2 summarizes the amplitudes $b$ and offsets $c$ of the relaxation traces for each stimulation duration $T$, which determine the functions $b(T)$ and $c(T)$ appearing in the cumulative conductance function-reconstruction procedure of IV.3. The cumulative conductance function power law coefficient is shown in Table˜3. The table˜4 lists cumulative conductance coefficients $h_{1}$ and $h_{2}$, as well as model coefficiets $\mu$ and $g_{0}$, estimated to fit the potentiation segment of Figure˜5 and Figure˜6.
Table˜5 reports the parameters of the STDP module (Sections˜II.2 and 3), which are used to generate the model STDP curve in Figure˜7. Our choice for the parameters of the saturation module are presented in table˜6
@@
Table 2: Numerical values of the trace-specific amplitudes $b$ and offsets $c$ obtained from fitting the relaxation traces for different stimulation durations $T$ (cf. IV.2). These values determine the functions $b(T)$ and $c(T)$ in Eq.˜10, which are used to derive the kernel function $\ker(t)$ and, in turn, the form of the cumulative conductance function $H(s)$ from the experimental data.
@@
3.5 s $1.70\times 10^{-6}$ $44.8\times 10^{-6}$ $0.226$ $3.07\times 10^{-5}$
35 s $2.78\times 10^{-6}$ $54.4\times 10^{-6}$ $0.215$ $3.08\times 10^{-5}$
@@
Table 5: Numerical values of the parameters of the STDP module used in Section˜IV.5. The parameters $_{\mathrm{+}}, aa_{\mathrm{-}}$ and $\tau_{\mathrm{+}}$, $\tau_{\mathrm{-}}$ define the shapes and time constants of the potentiating and depressing lobes in the local-variable STDP dynamics Eq.˜3, the coefficients $h_{\mathrm{01}}$ and $h_{\mathrm{00}}$ determine the dependence of the cumulative conductance function $H$ on the variable $w$ as in Eq.˜6. These values are used to generate the model STDP curve shown in Figure˜7 and to ensure consistency between the dynamical STDP formulation and the measured memristor response.
$g_{\mathrm{max}}$ $5\times 10^{-5}$
$g_{\mathrm{min}}$ $1.17\times 10^{-5}$

### 13 Apr 2026 v2

@@ Discussion
Extending the state-of-the-art memristor models
@@
To model synaptic plasticity, we implement a spike-timing-dependent plasticity (STDP) rule inspired by the work of Gerstner et al. [6]. This approach is based on the concept of eligibility traces, a cornerstone of modern theories of learning in the brain. In this framework, plasticity is a two-stage process. First, the near-coincidence of pre- and post-synaptic activity creates a short-lived synaptic "tag" or trace that marks the synapse as "eligible" for modification. Second, this temporary trace is only converted into a lasting weight change if a third, modulatory signal—representing global factors like reward, novelty, or surprise—arrives while the trace is still active [7]. By adopting this mechanism, our model moves beyond simple pairwise STDP to be extended toward and flexible three-factor learning rules. This capability aligns the model with cutting-edge research in reinforcement learning and makes it a compelling component for building truly adaptive neuromorphic systems.
@@
Physical origin of the 1/t-like memory kernel
The volatility module of our model is governed by the hereditary kernel in Eq.˜9 with an exponent $\alpha\approx 0.03$ obtained from the log–log relaxation data in Figure˜2. The corresponding cumulative memory $K(t)=\int_{-\infty}^{t}\ker(\tau)\,d\tau$ decays approximately as $K(t)\sim 1/t$ for large $t$. In practice, the device therefore exhibits effectively $1/t$-like “loss of memory” over time.
@@
The viscoelastic analogy provides a convenient heuristic explanation of this situation. In linear viscoelasticity, a material with a broad spectrum of relaxation times $\rho(\tau)$ can be represented as a superposition of many Maxwell modes, and a spectrum $\rho(\tau)\propto\tau^{-1-\alpha}$ yields an effective relaxation modulus $G(t)\propto t^{-\alpha}$ [15, 21]. Our choice of $\ker(t)\propto(t+\varepsilon)^{-(\alpha+1)}$ plays the same role: it compactly encodes the presence of many internal relaxation processes without having to resolve them individually at the level of the model.
@@
Adapting model to polymeric memristor
@@
Prospective model extensions
@@
Conclusions
We have presented a modular and computationally efficient compact model for memristors that exhibits both volatile memory and synaptic-like plasticity. The model’s key innovation lies in the synthesis of several modeling techniques that allow for memristor current-voltage characteristics, synaptic-like plasticity effects and a general description of relaxation dynamics with saturation. We have discussed how the volatile relaxation dynamics, experimentally observed as a power-law decay, can be described by a viscoelasticity-inspired hereditary integral whose form is physically motivated by percolation theory. Furthermore, we have demonstrated that a biologically plausible model of STDP based on eligibility traces can be implemented, endowing the future devices with a potential capacity for advanced learning.
The complete model was shown to be in excellent quantitative agreement with experimental data from polymeric memristors. Furthermore, we believe that the modularity and relative simplicity of the proposed model allow its implementation (including validation) in a wide range of memristive-like devices. Therefore, this work provides a unified, mathematically grounded framework for understanding and simulating complex memristive devices, offering a valuable tool for advancing the design of neuromorphic computing systems.
Funding
@@ Conclusions
Author contributions
D.H. and L.K. conceived and led the project, carried out the theoretical development, implemented the model and simulations, analyzed the data, and wrote the manuscript. S.H.F. and K.K. contributed to the theoretical development and provided critical feedback and revisions to specific sections of the manuscript. A.P., Y.R.P., J.P., and J.V. fabricated the polymeric memristors and performed the experimental measurements.
Competing interests
The authors declare no competing interests. All authors reviewed the manuscript and approved the final version.
@@
The source code for the model can be accessed at https://github.com/danielhabart/modular_memristor
Appendix: Numerical values
@@
The first Table˜1 lists the kernel exponent $\alpha$. The second table˜2 summarizes the amplitudes $b$ and offsets $c$ of the relaxation traces for each stimulation duration $T$, which determine the functions $b(T)$ and $c(T)$ appearing in the cumulative conductance function-reconstruction procedure Eq.˜10. The cumulative conductance function power law coefficient is shown in Table˜3. The Table˜4 lists cumulative conductance coefficients $h_{11}$ and $h_{10}$, as well as model coefficients $\mu$ and $g_{0}$, estimated to fit the potentiation segment of Figure˜5.
Table˜5 reports the parameters of the STDP module (Eq.˜3), which are used to generate the model STDP curve in Figure˜6. Our choice for the parameters of the saturation module are presented in Table˜6.
@@
Table 2: Numerical values of the trace-specific amplitudes $b$ and offsets $c$ obtained from fitting the relaxation traces for different stimulation durations $T$. These values determine the functions $b(T)$ and $c(T)$ in Eq.˜10, which are used to derive the kernel function $\ker(t)$ and, in turn, the form of the cumulative conductance function $H(s)$ from the experimental data.
@@
[…]
@@
Table 5: Numerical values of the parameters of the STDP module. The parameters $_{\mathrm{+}}, AA_{\mathrm{-}}$ and $\tau_{\mathrm{+}}$, $\tau_{\mathrm{-}}$ define the shapes and time constants of the potentiating and depressing lobes in the local-variable STDP dynamics Eq.˜3, the coefficients $h_{\mathrm{01}}$ and $h_{\mathrm{00}}$ determine the dependence of the cumulative conductance function $H$ on the variable $w$ as in Eq.˜6. These values are used to generate the model STDP curve shown in Figure˜6 and to ensure consistency between the dynamical STDP formulation and the measured memristor response.
[…]